# Time-tronics: from temporal printed circuit board to quantum computer

### v1

@@
Krzysztof Giergiel Instytut Fizyki Teoretycznej, Uniwersytet Jagielloński, ulica Profesora Stanislawa Lojasiewicza 11, PL-30-348 Kraków, Poland Optical Sciences Centre, Swinburne University of Technology, Melbourne 3122, Australia CSIRO, Manufacturing, Research Way, Clayton, Victoria 3168, Australia    Peter Hannaford Optical Sciences Centre, Swinburne University of Technology, Melbourne 3122, Australia    Krzysztof Sacha Instytut Fizyki Teoretycznej, Uniwersytet Jagielloński, ulica Profesora Stanislawa Lojasiewicza 11, PL-30-348 Kraków, Poland
@@
Time crystalline structures can be created in periodically driven systems (Sacha, 2020; Zaletel et al., 2023). They are temporal lattices which can reveal different condensed matter behaviours ranging from Anderson localization in time to temporal analogues of many-body localization or topological insulators Sacha (2020); Hannaford and Sacha (2022). However, the potential practical applications of time crystalline structures have yet to be explored. Here, we pave the way for time-tronics where temporal lattices are like printed circuit boards for realization of a broad range of quantum devices. The elements of these devices can correspond to structures of dimensions higher than three and can be arbitrarily connected and reconfigured at any moment. Moreover, our approach allows for the construction of a quantum computer, enabling quantum gate operations for all possible pairs of qubits. Our findings indicate that the limitations faced in building devices using conventional spatial crystals can be overcome by adopting crystalline structures in time.
@@
Time crystals, much like their spatial counterparts, spontaneously form through the self-organization of many-body systems but in the time domain Wilczek (2012); Bruno (2013); Watanabe and Oshikawa (2015); Kozin and Kyriienko (2019). In periodically driven systems, discrete time crystals can emerge, evolving spontaneously with a period longer than that dictated by the periodic perturbation Sacha (2015a); Khemani et al. (2016); Else et al. (2016); Zhang et al. (2017); Choi et al. (2017); Pal et al. (2018); Rovny et al. (2018); Smits et al. (2018); Mi et al. (2022); Randall et al. (2021); Frey and Rachel (2022); Keßler et al. (2021); Kyprianidis et al. (2021); Xu et al. (2021); Taheri et al. (2022); Bao et al. (2024); Shinjo et al. (2024); Liu et al. (2024a, b). In periodically disturbed systems, it is also possible to create time lattices, not resulting from self-organization, but from the application of appropriate temporal resonant perturbations Guo et al. (2013); Sacha (2015b); Buchleitner et al. (2002). Studies have shown that such time crystalline structures can manifest a wide range of phases known in the physics of condensed matter, such as Anderson localization and many-body localization in time, topological insulators in time, and Mott insulators in the time domain Sacha (2020); Guo (2021); Hannaford and Sacha (2022).
@@
From the inception of the field of time crystals, questions have arisen regarding how crystalline structures in time can be practically utilized and whether they carry greater potential for practical applications than conventional spatial crystals Sacha (2020); Hannaford and Sacha (2022); Zaletel et al. (2023); Bomantara and Gong (2018); Estarellas et al. (2020); Carollo et al. (2020); Lyu et al. (2020); Iemini et al. (2023); Montenegro et al. (2023); Bao et al. (2024); Yousefjani et al. (2024). In this article, we demonstrate that it is possible to achieve time-tronics, where time crystalline structures serve as printed circuit boards as in electronics, allowing the design and realization of a broad range of quantum devices.
@@
Next, we focus on a specific example, a quantum computer Preskill (2018). We demonstrate that a temporal printed circuit board can host qubits, where all single-qubit operations can be realized and a controlled-Z gate can be performed between all possible qubit pairs, meeting the conditions for a universal quantum computer.
Temporal printed circuit board
@@
To explain the creation of a time crystalline structure, let us begin with a classical particle moving periodically in a one-dimensional (1D) box between infinite walls. Such a periodic trajectory remains stable in the presence of a periodically oscillating potential if the resonance condition is met, i.e., if the frequency $\omega$ of the oscillating potential is $s$ times greater than the frequency of the particle motion, where $s$ is an integer number Buchleitner et al. (2002). There are $s$ oscillations of the potential for each period of the particle trajectory, allowing for the simultaneous driving of $s$ particles that follow each other with a delay $T=2\pi/\omega$, as illustrated in Fig. 1a (see also Methods).
@@
Figure 1: (a): Red ball illustrates a classical particle moving in a 1D box potential in the presence of an oscillating spatially periodic potential. If the period of particle motion is $s$ times longer than the period of the oscillating potential, $T=2\pi/\omega$, we have an $s:1$ resonance. In such a case, $s$ classical non-interacting particles can be positioned so that they follow one another along the resonant trajectory. In the quantum description, the $s:1$ resonance manifests itself with the presence of $s$ localized wave-packets, which evolve one after another along the resonant trajectory. A quantum particle prepared in a certain wave-packet can tunnel to neighboring wave-packets. Here, we assume that the wave-packets are sufficiently localized so that natural tunneling is negligible. (b): Tunneling can be established in a controlled manner using Bragg scattering Müller et al. (2008). If during the encounter of the $i$-th and $j$-th wave-packets moving in opposite directions, we briefly turn on one broad laser beam (labelled $\mathbf{k}_{2}$) and one narrow laser beam (labelled $\mathbf{k}_{1}$) focused on the meeting point of the wave-packets and the Bragg condition is satisfied, we can realize atom tunneling between the wave-packets with an amplitude $J_{ij}$, the modulus of which depends on the beam parameters and the interaction time with the atom, and a phase depending on the relative phase between the beams. With a controllable array of focused laser beams available (e.g., using a digital micromirror device Wang et al. (2020) as sketched in the figure), which can be independently activated at different time points, one can control the tunneling amplitudes $J_{ij}$ between any pair of wave-packets. (c): Ultra-cold atoms can be prepared in states where they do not interact. In such a situation, we have the possibility of selectively controlling interactions between atoms occupying any pair of wave-packets. If just before the encounter of atoms occupying two wave-packets, we perform a Raman transfer Levine et al. (2022) (using broad and focused laser beams) from the initial internal states of the atoms to states where the atoms interact, then during the passage of the wave-packets, interaction between atoms will occur. After the wave-packets pass each other, we perform a Raman transfer again, but back to non-interacting internal states of the atoms. With a digital micromirror device available, interaction between atoms occupying any pair of wave-packets can be realized. (d): Error of atom transfer between two wave-packets using Bragg pulses versus the assumed number of cycles of the resonant trajectory needed to achieve full transfer (two Bragg pulses per each cycle). Faster transfer requires stronger Bragg pulses, $\lambda_{\rm Bragg}$, increasing the coupling of the atom to other undesired states. Generally, apart from additional coherent oscillations, the longer the realization of the transfer, the smaller the error. (e): Fidelity of the controlled-Z gate (CZ), where the interaction imparts a phase of $\pi$ to the state where one atom occupies one wave-packet and the second atom occupies another wave-packet, versus the number of required cycles (two interaction meetings of the wave-packets per cycle). The two curves represent the cases without (black) and with (red) the inclusion of Raman transfer errors Bluvstein et al. (2022) (see Methods). For a small number of cycles, stronger interactions are needed, and the coupling to other states limits the CZ fidelity. For a larger number of cycles, more Raman transfers are needed, and their imperfections limit the fidelity. The results presented in (d) and (e) correspond to 39K atoms driven resonantly by an optical lattice potential, created by laser radiation with a wavelength $10.6\;\mu$m, which oscillates with a frequency 5.46 kHz and with Bragg and Raman pulses of wavelength 266 nm (see Methods).
@@
In the quantum realm, this classical system has its close counterpart, where each particle is represented by a localized wave-packet propagating without spreading along the resonant trajectory with a period $sT$ Buchleitner et al. (2002); Sacha (2020). In the quantum world, there is the tunneling process that can lead to the transfer of a particle from one wave-packet to neighboring wave-packets. Tunneling times become very long and thus negligible in the experiment if the wave-packets are sufficiently strongly localized. In the following, we assume negligible natural tunneling, but we can selectively introduce tunneling between any wave-packets in a controlled manner using Bragg scattering Müller et al. (2008), as explained in Fig. 1b (see also Methods). In summary, this allows us to create a system of non-interacting bosons (or fermions) described by the tight-binding model,
@@
Initially, we assume non-interacting atoms. This scenario can be realized by means of a Feshbach resonance, i.e., an external magnetic field can be chosen so that the s-wave scattering length which describes the contact interactions between atoms is zero Pethick and Smith (2002). The interactions between atoms can then be selectively reinstated, even between atoms occupying different lattice sites in (1), i.e., with $i\neq j$. A similar array of laser beams as used for selective Bragg scattering and control of the tunneling amplitudes $J_{ij}$ can serve this purpose. These laser beams allow for a selective change in the internal state of the atoms by Raman transfer occupying given wave-packets Levine et al. (2022). This change leads to interaction between the atoms when their respective wave-packets intersect during the periodic evolution along the resonant trajectory because the scattering length is no longer zero, as depicted in Fig. 1c. Subsequent to the interaction, the atoms can be reverted to their initial internal state. This approach allows for precise control over interactions between atoms occupying any pair of wave-packets which are characterized by the coefficients $U_{ij}$ in the Hamiltonian (see Methods)
@@
Figure 2: Temporal printed circuit board. (a): The $s$ wave-packets evolving along the resonant trajectory (Fig. 1a) can be treated as $s$ states of a lattice with $s$ sites. If we arrange the sites of the lattice in a 1D chain and using Bragg scattering (Fig. 1b) we realize tunneling between nearest neighbors, we will have a 1D crystalline structure with nearest neighbor hoppings. If we arrange the sites in a 2D lattice and tunneling between nearest neighbors, we will have a 2D crystalline structure. Similarly, we can realize 3D and higher-dimensional crystalline structures. (b): Let us consider $s$ periodically evolving wave-packets as states of a 2D lattice. Using a broad laser beam and having only two focused laser beams at our disposal (Fig. 1b), all nearest neighbor tunnelings of the 2D lattice can be realized. For example, a 2D sublattice can be coupled to another 2D sublattice via a 1D chain as shown in the figure (realized tunnelings are marked with blue arrows). Control over the phases of tunneling amplitudes $J_{ij}$ allows for the generation of an artificial magnetic field in the system. With an array of focused laser beams available, any sites in the lattice can be connected via tunneling, as illustrated by the long blue arrow. More exotic geometries can also be realized, such as the Klein bottle, where the left and right edges of the 2D lattice are connected normally, but the top edge is twisted before connection to the bottom edge. Using selective Raman transfer (Fig. 1c), interactions between atoms occupying any pair of sites can be realized and controlled (as indicated by red arched arrows in the plot). During an experiment, the entire system can undergo any reconfiguration.
@@
With the universal time crystalline structure described by the Hamiltonian (2), it is possible to realize a broad range of quantum devices. Figure 2a presents a few examples. Physically the states associated with the sites of the time crystalline lattice are localized wave-packets moving periodically in a 1D box (Fig. 1a). One can represent the crystalline structure as a row of sites but if the sites are connected via selective Bragg tunneling as depicted in the middle plot in Fig. 2a, then the structure can be represented as a 2D lattice with nearest neighbor tunnelings, and 2D condensed matter problems can be realized and investigated. For example, fermionic or bosonic atoms can experience a magnetic-like field if proper complex phases of the tunneling amplitudes $J_{ij}$ are realized by a proper choice of the relative phase of the laser beams in the Bragg scattering (see Methods). A few such 2D lattices can be created and connected via selective tunneling to form a 3D lattice (bottom plot in Fig. 2a), and this procedure can be continued to form a 4D lattice, and so on. It is also possible to realize exotic objects like a Klein bottle, which is hard to imagine in 3D space. That is, pairs of opposite edges of a 2D lattice can be connected via tunneling, but one of the edges is twisted before the connection.
@@
Figure 3: Quantum computer. (a): Initially, $s/2$ bosonic atoms are prepared in a Mott insulator phase in a 1D box potential in the presence of a static optical lattice. We assume there are $s/2$ potential wells in the box, with one atom in each well. After preparing the Mott insulator phase, the interactions between atoms are turned off by means of a Feshbach resonance Pethick and Smith (2002). The next step is to impart a kick to the atoms, so that they start moving with momenta that satisfy the $s:1$ resonance condition with the frequency of the optical lattice oscillation, which we simultaneously switch on. Shortly after the previously described procedure, we observe $s/2$ wave-packets occupied by single atoms moving to the right in the figure and $s/2$ unoccupied wave-packets moving to the left. We assign one unoccupied wave-packet to each occupied wave-packet, forming the $|0\rangle$ and $|1\rangle$ states of qubits. In total, we have $s/2$ qubits. When two wave-packets corresponding to the $|0\rangle$ and $|1\rangle$ states of the same qubit meet during evolution along the resonant trajectory, a single-qubit gate can be performed using Bragg scattering. Control over the relative phase of the laser beams used in the Bragg scattering allows us to control whether the $\sigma_{x}$ or $\sigma_{y}$ operation is performed. High fidelity single-qubit operations are achieved by dividing them into several stages, meaning the entire single operation requires multiple encounters of wave-packets, and hence spanning several periods of the resonant trajectory (Fig. 1d). (b): When wave-packets corresponding to the states $|1\rangle_{i}$ and $|1\rangle_{j}$ of the $i$-th and $j$-th qubits meet during evolution, a controlled-Z gate can be realized. Atoms initially do not interact, but just before the wave-packets pass each other, we change the internal states of the atoms to states where they interact. We do this using Raman transfer with two beams. After passing each other, another Raman transfer restores the initial internal states of the atoms. If the interaction strength between atoms and the duration of interaction are appropriately chosen, after the wave-packets pass, the state $|1\rangle_{i}|1\rangle_{j}$ acquires a phase $e^{i\pi}$, and the controlled-Z gate is completed. Similarly to single-qubit operations, a high fidelity controlled-Z gate can be realized by dividing it into several encounters of the appropriate wave-packets (Fig. 1e). (c): Quantum volume Cross et al. (2018) calculated based on the B-gate decomposition of a generic SU(4) two-qubit operation Zhang et al. (2003). In this plot we assume no error in the single qubit operations. We plot a square of fidelilty of decomposition of a single B-gate using optimal points selected from the CZ-fidelity plot (Fig. 1e). For the higher the Raman transfer fidelity one can use a longer multi-cycle realization of the CZ-gates leading to quick increase in the quantum volume. The results correspond to the same parameters as in Figs. 1d-1e.
Quantum computer
@@
The construction of a universal quantum computer requires the preparation of numerous qubits (e.g., two-level atoms), the ability to perform all single-qubit operations, and the capability to execute two-qubit operations (e.g., controlled-Z gate operation) between any pairs of qubits Preskill (2018). The latter necessitates selective interactions between atoms, which, in turn, demand precise coherent transport of atoms and activation of interactions when they are in close proximity Cirac and Zoller (2000); Bluvstein et al. (2022); González-Cuadra et al. (2023).
@@
Crystalline structures in time automatically address the atom transport problem, as atoms are prepared in wave-packets evolving along a periodic trajectory, and each atom individually encounters every other atom at some point in time. Additionally, every atom moving on the same trajectory means they experience the same low frequency spatially inhomogenous fields, thus likely reducing phase noise sources. To realize a controlled-Z gate for any pair of atoms, it is sufficient to activate interactions between them at the moment of their encounter. The temporal printed circuit board described in the previous section is well-suited for this purpose. The board also enables the implementation of all necessary single-qubit operations required for building a quantum computer. For the successful construction of a quantum computer, all gate operations need to be as straightforward as possible. In the proposed approach, which we describe in the following, we only need a static arrangement of focused laser beams that will be activated at the appropriate moments, allowing the implementation of all necessary gate operations.
@@
The crystalline structure in time, as described in Fig. 1, consists of $s$ states ($s$ localized wave-packets). In such a structure, if we prepare $s/2$ atoms, we can define $s/2$ qubits, assigning two different wave-packets to each atom (see Fig. 3). If the atom occupies the first chosen wave-packet, the qubit will be in the state $|1\rangle$ and if it occupies the second wave-packet, it will be in the state $|0\rangle$. The initialization process of all qubits in, for example, state $|1\rangle$ is relatively straightforward. In the 1D box potential and in the presence of a static deep optical lattice, we prepare bosons in a Mott insulator state with unit filling of lattice sites Pethick and Smith (2002), assuming the number of lattice sites between the walls of the box potential is $s/2$ (see Methods). Next, we turn off the interactions between atoms setting the s-wave scattering to zero by means of a Feshbach resonance Pethick and Smith (2002). We then kick the atoms so that they start moving in one direction with momentum $p$ that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation and activate the optical lattice oscillation. This leads to a situation where $s/2$ wave-packets of the time crystalline structure that are initially moving in the same direction are occupied by individual atoms, while the other $s/2$ wave-packets initially moving in the opposite direction are unoccupied, see Fig. 3a and Methods. For each occupied wave-packet, we assign any unoccupied wave-packet to form a qubit together. The choice of which unoccupied wave-packets are assigned to which occupied wave-packets is arbitrary.

### v2

@@
[…]
@@
Time crystalline structures can be created in periodically driven systems [1, 2]. They are temporal lattices which can reveal different condensed matter behaviours ranging from Anderson localization in time to temporal analogues of many-body localization or topological insulators [1, 3]. However, the potential practical applications of time crystalline structures have yet to be explored. Here, we pave the way for time-tronics where temporal lattices are like printed circuit boards for realization of a broad range of quantum devices. The elements of these devices can correspond to structures of dimensions higher than three and can be arbitrarily connected and reconfigured at any moment. Moreover, our approach allows for the construction of a quantum computer, enabling quantum gate operations for all possible pairs of qubits. Our findings indicate that the limitations faced in building devices using conventional spatial crystals can be overcome by adopting crystalline structures in time.
@@
Time crystals, much like their spatial counterparts, spontaneously form through the self-organization of many-body systems but in the time domain [4, 5, 6, 7]. In periodically driven systems, discrete time crystals can emerge, evolving spontaneously with a period longer than that dictated by the periodic perturbation [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In periodically disturbed systems, it is also possible to create time lattices, not resulting from self-organization, but from the application of appropriate temporal resonant perturbations [27, 28, 29]. Studies have shown that such time crystalline structures can manifest a wide range of phases known in the physics of condensed matter, such as Anderson localization and many-body localization in time, topological insulators in time, and Mott insulators in the time domain [1, 30, 3].
@@
From the inception of the field of time crystals, questions have arisen regarding how crystalline structures in time can be practically utilized and whether they carry greater potential for practical applications than conventional spatial crystals [1, 31, 2, 32, 33, 34, 35, 36, 37, 23, 38]. In this article, we demonstrate that it is possible to achieve time-tronics, where time crystalline structures serve as printed circuit boards as in electronics, allowing the design and realization of a broad range of quantum devices.
@@
Next, we focus on a specific example, a quantum computer [39]. We demonstrate that a temporal printed circuit board can host qubits, where all single-qubit operations can be realized and a controlled-Z gate can be performed between all possible qubit pairs, meeting the conditions for a universal quantum computer.
@@
Temporal printed circuit board. To explain the creation of a time crystalline structure, let us begin with a classical particle moving periodically in a one-dimensional (1D) box between infinite walls. Such a periodic trajectory remains stable in the presence of a periodically oscillating potential if the resonance condition is met, i.e., if the frequency $\omega$ of the oscillating potential is $s$ times greater than the frequency of the particle motion, where $s$ is an integer number [29]. There are $s$ oscillations of the potential for each period of the particle trajectory, allowing for the simultaneous driving of $s$ particles that follow each other with a delay $T=2\pi/\omega$, as illustrated in Fig. 1a (see also Methods).
@@
Figure 1: (a): Red ball illustrates a classical particle moving in a 1D box potential in the presence of an oscillating spatially periodic potential. If the period of particle motion is $s$ times longer than the period of the oscillating potential, $T=2\pi/\omega$, we have an $s:1$ resonance. In such a case, $s$ classical non-interacting particles can be positioned so that they follow one another along the resonant trajectory. In the quantum description, the $s:1$ resonance manifests itself as $s$ localized wave-packets, which evolve one after another along the resonant trajectory. Here, we assume that the wave-packets are sufficiently localized so that natural tunneling between neighboring wave-packets is negligible. (b): If during the encounter of the $i$-th and $j$-th wave-packets moving in opposite directions, one broad laser beam (labelled $\mathbf{k}_{2}$) and one narrow laser beam (labelled $\mathbf{k}_{1}$) focused on the meeting point of the wave-packets are briefly turned on and the Bragg scattering [40] condition is satisfied, we can realize atom tunneling between the wave-packets with an amplitude $J_{ij}$, the modulus of which depends on the beam parameters and the interaction time with the atom, and a phase depending on the relative phase between the beams. With a controllable array of focused laser beams available (e.g., using a digital micromirror device [41] as in the figure), which can be independently activated at different time points, one can control the tunneling amplitudes $J_{ij}$ between any pair of wave-packets. (c): Ultra-cold atoms can be prepared via a Feshbach resonance in states where they do not interact. In such a situation, we have the possibility of selectively controlling interactions between atoms occupying any pair of wave-packets. If just before the encounter of atoms occupying two wave-packets, we perform a Raman transfer [42] (using broad and focused laser beams) from the initial internal states of the atoms to states where the atoms interact, then during the passage of the wave-packets, interaction between atoms will occur. After the wave-packets pass each other, we perform a Raman transfer again, but back to non-interacting internal states of the atoms. With a digital micromirror device available, interaction between atoms occupying any pair of wave-packets can be realized. (d): Error of atom transfer between two wave-packets using Bragg pulses versus the assumed number of cycles of the resonant trajectory needed to achieve full transfer (two Bragg pulses per each cycle). Faster transfer requires stronger Bragg pulses, $\lambda_{\rm Bragg}$, increasing the coupling of the atom to other undesired states. Generally, apart from additional coherent oscillations, the longer the realization of the transfer, the smaller the error. (e) and (f): Fidelity of the controlled-Z gate (CZ), in which the interaction imparts a phase of $\pi$ to the state where one atom occupies one wave-packet and the second atom occupies another wave-packet, versus the number of required cycles (two interaction meetings of the wave-packets per cycle). The two curves represent the cases without (black) and with (red) the inclusion of Raman transfer errors [43] (see Methods). For a small number of cycles, stronger interactions are needed, and the coupling to other states limits the CZ fidelity. For a larger number of cycles, more Raman transfers are needed, and their imperfections limit the fidelity. (e) is for a variable interaction $g_{0}$ and fixed Raman transfer time ($T_{\rm Raman}\approx 0.12T$, before and after the center-point meeting time) and (f) is for a fixed interaction ($g_{0}=10$) and variable Raman transfer time. The results presented in (d)-(f) correspond to 39K atoms driven resonantly by an optical lattice potential, created by laser radiation with a wavelength $10.6\;\mu$m, which oscillates with a frequency 5.46 kHz and with Bragg pulses of wavelength 266 nm (see Methods).
@@
In the quantum realm, this classical system has its close counterpart, where each particle is represented by a localized wave-packet propagating without spreading along the resonant trajectory with a period $sT$ [29, 1]. In the quantum world, there is the tunneling process that can lead to the transfer of a particle from one wave-packet to neighboring wave-packets. Tunneling times become very long and thus negligible in the experiment if the wave-packets are sufficiently strongly localized. In the following, we consider ultra-cold atoms and assume negligible natural tunneling, but we can selectively introduce tunneling between any wave-packets in a controlled manner using Bragg scattering [40], as explained in Fig. 1b (see also Methods). In summary, this allows us to create a system of non-interacting bosons (or fermions) described by the tight-binding model,
@@
Initially, we assume non-interacting atoms. This scenario can be realized by means of a Feshbach resonance, i.e., an external magnetic field can be chosen so that the s-wave scattering length which describes the contact interactions between atoms is zero [44]. The interactions between atoms can then be selectively reinstated, even between atoms occupying different lattice sites in (1), i.e., with $i\neq j$. A similar array of laser beams as used for selective Bragg scattering and control of the tunneling amplitudes $J_{ij}$ can serve this purpose. These laser beams allow for a selective change in the internal state of the atoms by Raman transfer occupying given wave-packets [42]. This change leads to interaction between the atoms when their respective wave-packets intersect during the periodic evolution along the resonant trajectory because the scattering length is no longer zero, as depicted in Fig. 1c. Subsequent to the interaction, the atoms can be reverted to their initial internal state. This approach allows for precise control over interactions between atoms occupying any pair of wave-packets which are characterized by the coefficients $U_{ij}$ in the Hamiltonian (see Methods)
@@
Figure 2: Temporal printed circuit board. (a): The $s$ wave-packets evolving along the resonant trajectory (Fig. 1a) can be treated as $s$ states of a lattice with $s$ sites. If we arrange the sites of the lattice in a 1D chain and using Bragg scattering (Fig. 1b) we realize tunneling between nearest neighbors, we will have a 1D crystalline structure with nearest neighbor hoppings. If we arrange the sites in a 2D lattice and induce tunneling between nearest neighbors, we will have a 2D crystalline structure. Similarly, we can realize 3D and higher-dimensional crystalline structures. (b): Let us consider $s$ periodically evolving wave-packets as states of a 2D lattice. Using a broad laser beam and having only two focused laser beams at our disposal (Fig. 1b), all nearest neighbor tunnelings of the 2D lattice can be realized. For example, a 2D sublattice can be coupled to another 2D sublattice via a 1D chain as shown in the figure (realized tunnelings are marked with short blue arrows). Control over the phases of tunneling amplitudes $J_{ij}$ allows for the generation of an artificial magnetic field in the system. With an array of focused laser beams available, any sites in the lattice can be connected via tunneling, as illustrated by the long blue arrow. More exotic geometries can also be realized, such as the Klein bottle, where the left and right edges of the 2D lattice are connected normally, but the top edge is twisted before connection to the bottom edge. Using selective Raman transfer (Fig. 1c), interactions between atoms occupying any pair of sites can be realized and controlled (as indicated by red arched arrows in the plot). During an experiment, the entire system can undergo any reconfiguration.
@@
With the universal time crystalline structure described by the Hamiltonian (2), it is possible to realize a broad range of quantum devices. Figure 2a presents a few examples. Physically, the states associated with the sites of the time crystalline lattice are localized wave-packets moving periodically in a 1D box (Fig. 1a). One can represent the crystalline structure as a row of sites but if the sites are connected via selective Bragg tunneling as depicted in the middle plot in Fig. 2a, then the structure can be represented as a 2D lattice with nearest neighbor tunnelings, and 2D condensed matter problems can be realized and investigated. For example, fermionic or bosonic atoms can experience a magnetic-like field if proper complex phases of the tunneling amplitudes $J_{ij}$ are realized by a proper choice of the relative phase of the laser beams in the Bragg scattering (see Methods). A few such 2D lattices can be created and connected via selective tunneling to form a 3D lattice (bottom plot in Fig. 2a), and this procedure can be continued to form a 4D lattice, and so on. It is also possible to realize exotic objects like a Klein bottle, which is hard to imagine in 3D space [45, 46]. That is, pairs of opposite edges of a 2D lattice can be connected via tunneling, but one of the edges is twisted before the connection.
@@
Figure 3: Quantum computer. (a): Initially, $s/2$ bosonic atoms are prepared in a Mott insulator phase in a 1D box potential in the presence of a static optical lattice. We assume there are $s/2$ potential wells in the box, with one atom in each well. After preparing the Mott insulator phase, the interactions between atoms are turned off by means of a Feshbach resonance [44]. The next step is to impart momentum to the atoms that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation, which we simultaneously switch on. Shortly after the previously described procedure, we observe $s/2$ wave-packets occupied by single atoms moving to the right in the figure and $s/2$ unoccupied wave-packets moving to the left. We assign one unoccupied wave-packet to each occupied wave-packet, forming the $|0\rangle$ and $|1\rangle$ states of qubits. In total, we have $s/2$ qubits. When two wave-packets corresponding to the $|0\rangle$ and $|1\rangle$ states of the same qubit meet during evolution along the resonant trajectory, a single-qubit gate can be performed using Bragg scattering. Control over the relative phase of the laser beams used in the Bragg scattering allows us to control whether the $\sigma_{x}$ or $\sigma_{y}$ operation is performed. High fidelity single-qubit operations are achieved by dividing them into several stages, meaning the entire single operation requires multiple encounters of wave-packets, and hence spanning several periods of the resonant trajectory (Fig. 1d). (b): When wave-packets corresponding to the states $|1\rangle_{i}$ and $|1\rangle_{j}$ of the $i$-th and $j$-th qubits meet during evolution, a controlled-Z gate can be realized. Atoms initially do not interact, but just before the wave-packets pass each other, we change the internal states of the atoms to states where they interact. We do this using Raman transfer with two beams. After the wave-packets pass each other, another Raman transfer restores the initial internal states of the atoms. If the interaction strength between atoms and the duration of interaction are appropriately chosen, after the wave-packets pass, the state $|1\rangle_{i}|1\rangle_{j}$ acquires a phase $e^{i\pi}$, and the controlled-Z gate is completed. Similarly to single-qubit operations, a high fidelity controlled-Z gate can be realized by dividing it into several encounters of the appropriate wave-packets (Fig. 1e-1f). (c): Quantum volume [47] calculated based on the B-gate decomposition of a generic SU(4) two-qubit operation [48]. In this plot we assume no error in the single-qubit operations. We plot a square of fidelilty of decomposition of a single B-gate using optimal points selected from the CZ-fidelity plot (Fig. 1e). For higher Raman transfer fidelity one can use a longer multi-cycle realization of the CZ-gates leading to quick increase in the quantum volume. The results correspond to the same parameters as in Figs. 1d-1f.
@@
Quantum computer. As an important example of the application of a temporal printed circuit board, we consider the case of a quantum computer. The construction of a universal quantum computer requires the preparation of numerous qubits (e.g., two-level atoms), the ability to perform all single-qubit operations, and the capability to execute two-qubit operations (e.g., controlled-Z gate operation) between any pairs of qubits [39]. The latter necessitates selective interactions between atoms, which, in turn, demand precise coherent transport of atoms and activation of interactions when they are in close proximity [49, 43, 50].
@@
Crystalline structures in time automatically address the atom transport problem, as atoms are prepared in wave-packets evolving along a periodic trajectory, and each atom individually encounters every other atom at some point in time. Additionally, every atom moving on the same trajectory means they experience the same spatially inhomogenous fields, thus likely reducing sources of phase noise. To realize a controlled-Z gate for any pair of atoms, it is sufficient to activate interactions between them at the moment of their encounter. The temporal printed circuit board described in the previous section is well-suited for this purpose. The board also enables the implementation of all necessary single-qubit operations required for building a quantum computer. For the successful construction of a quantum computer, all gate operations need to be as straightforward as possible. In the proposed approach, which we describe in the following, we only need a static arrangement of focused laser beams that will be activated at the appropriate moments, allowing the implementation of all necessary gate operations.
@@
The crystalline structure in time, as described in Fig. 1, consists of $s$ states ($s$ localized wave-packets). In such a structure, if we prepare $s/2$ atoms, we can define $s/2$ qubits, assigning two different wave-packets to each atom (see Fig. 3). If the atom occupies the first chosen wave-packet, the qubit will be in the state $|1\rangle$ and if it occupies the second wave-packet, it will be in the state $|0\rangle$. The initialization process of all qubits in, for example, state $|1\rangle$ is relatively straightforward. In the 1D box potential and in the presence of a static deep optical lattice, we prepare bosons in a Mott insulator state with unit filling of lattice sites [44], assuming the number of lattice sites between the walls of the box potential is $s/2$ (see Methods). Next, we turn off the interactions between atoms setting the s-wave scattering length to zero by means of a Feshbach resonance [44]. We then set the atoms moving in one direction with momentum $p$ that satisfies the $s:1$ resonance condition with the frequency of the optical lattice oscillation. To do this, we set the initially static optical lattice in motion so that in the moving frame the atoms follow the ground state of the moving potential well. This is achieved using a counter-diabatic driving process which involves a unitary transformation to the moving frame by applying a linear potential proportional to the acceleration with which the lattice is set in motion (see Supplemental Material for details). We then activate the optical lattice oscillation. This leads to a situation where $s/2$ wave-packets of the time crystalline structure that are initially moving in the same direction are occupied by individual atoms, while the other $s/2$ wave-packets initially moving in the opposite direction are unoccupied, see Fig. 3a and Methods. For each occupied wave-packet, we assign any unoccupied wave-packet to form a qubit together. The choice of which unoccupied wave-packets are assigned to which occupied wave-packets is arbitrary.